# A Statistician’s Overview of Mechanistic-Informed Modeling

### v1

@@ 3.2.2 Probabilistic Numerical Methods
Over the last decade, there has been growing interest in modernizing the ideas associated with treating differential equation solutions as random processes (e.g., Larkin,, 1972; Diaconis,, 1988). These ideas, which come under a broad set of methods called probabilistic numerical methods (PNMs), have provided a comprehensive toolkit to consider uncertainty quantification in the solution of differential equations (e.g., see Hennig et al.,, 2015; Chkrebtii et al.,, 2016; Cockayne et al.,, 2019, and references therein). These methods have the advantage of being quite rigorous in their convergence guarantees and are straightforward to implement.
Many PNM methods for solving PDEs with uncertainty assume the solution has a GP prior. Then observations of boundary conditions and forcing terms of the PDE at a finite set of points act as data that allow for a posterior distribution over the PDE solution. For a simple illustration, in the space-only case, we might have the following GP prior:
@@
$f(s_{i})=\mathcal{M}(u(s_{i}))+\epsilon_{i},\;\;\epsilon_{i}\sim\text{Gau}(0,\sigma^{2}_{\epsilon}),$
where $\sigma^{2}_{\epsilon}$ controls the fidelity of the PDE solution. Given the prior and data models are Gaussian, the posterior distribution of $u(s)$ given $f(s_{1}),\ldots,f(s_{m})$ are available in closed form via a Gaussian distribution. In practice, one includes a boundary condition process and, for time-varying PDEs, an initial condition. In addition, one typically chooses “collocation” points for which to obtain “observations” $f(s_{i})$ to facilitate computation. This approach is often embedded in a BHM with real world observations to learn the parameters of the underlying PDE (i.e., the inverse problem).
The strength of the PNM approach is its principled uncertainty quantification, which comes from the numerical approximation (e.g., the $f(s_{i})$ observations) and, in the case of inverse problems, real-world observations. The inherent regularization of the Bayesian prior also provides some robustness to ill-posedness. Perhaps most importantly, PNM’s strong connections to traditional numerical methods have been established for special cases (e.g., linear PDEs and GP priors). That said, these methods are much more difficult to implement for nonlinear PDEs, do not scale well for very high-dimensional problems, and have not been used as much in deep BHMs with multiple sources of process uncertainty, as in the physical-statistical paradigm. However, the Bayesian PINN framework described in the next section can be considered a neural network based PNM.
@@ 4 Physics-Informed Neural Networks (PINNs)
In the spirit of the Sasaki’s variational approach described above in Section 3.1, consider fitting a deep neural network but constraining the solution to follow a PDE. This is the essence of PINNs (Raissi et al., 2017a, ; Raissi et al., 2017b, ; Raissi,, 2018; Raissi et al.,, 2019, 2020). Such approaches can be used in the context of fitting a PDE in the presence of data (as in data assimilation), or solving a PDE explicitly (presumably for the sake of efficiency). We are interested here in the situation where we have data corresponding to the latent process of interest, initial conditions, or boundary conditions. We provide a brief overview of the PINN approach and its Bayesian counterpart (B-PINN) in this section, and then describe how these can fit into a mechanistically-informed BHM framework in Section 5.
@@ 4.1 PINN Formulation
PINNs, first proposed by Raissi et al., 2017a ; Raissi et al., 2017b , mainly consist of two components: (1) a neural network for prediction and (2) a physically-inspired equation (e.g., PDE) to impose constraints in the loss function. We denote the deep neural network solution to the PDE at time $t$ and location $\bm{s}$ as $\bm{u}_{NN,t}(\bm{s};\bm{\theta}_{W})$, where $\bm{\theta}_{W}$ represents the neural network parameters (weights and biases). Note, $\bm{u}_{NN}$ is an approximation of the PDE solution $\tilde{\bm{u}}$. Indeed, since this is not a perfect representation it is best to make the distinction as follows by defining a PDE “residual” (as in the PNM approach described in Section 3.2.2). Let $\bm{\theta}_{m}$ correspond to PDE parameters as in (1), and write
$\bm{r}_{t}(\bm{s};\bm{\theta}_{W},\bm{\theta}_{m})\equiv\frac{\partial{\bm{u}_{NN}(\bm{\theta}_{W})}}{\partial{t}}-\mathcal{M}_{u}({\bm{u}_{NN}(\bm{\theta}_{W})};\bm{\theta}_{m}),$ (8)
@@
$J_{IC}(\bm{\theta}_{W})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}\|{\bm{u}}_{NN,0}(\bm{s}_{IC}^{i};\bm{\theta}_{W})-\bm{g}_{I}(\bm{s}_{IC}^{i})\|_{2}^{2},$
$J_{BC}(\bm{\theta}_{W})=\frac{1}{N_{BC}}\sum_{i=1}^{N_{BC}}\|\bm{B}({\bm{u}}_{NN,t_{BC}^{i}}(\bm{s}_{BC}^{i};\bm{\theta}_{W}))\|_{2}^{2}.$
@@
where the specified non-negative constants $\{a_{IC},a_{BC},a_{P}\}$ allow one to focus more on different components of the solution if desired. The solution in this case does not require measurements of the process besides initial and boundary condition training data, so the PINN serves as an alternative to the traditional numerical solver (e.g., finite difference, finite element or spectral scheme). In this context, note a well-known limitation of the PINN approach is it can have difficulty in simulating multi-scale, chaotic, or turbulent behavior because the explicit forward (“causal”) nature of time is not considered (e.g., Wang et al.,, 2022). This limitation can be mediated in a number of ways, for example, by adding a time weighting function in the loss function (9). Regardless, some recent comparison studies have shown that traditional finite element solutions of PDEs work better and are as or more efficient than PINNs for most problems (e.g., Grossmann et al.,, 2024).
@@ 4.1.1 PINNs with Observational Data
$J_{D}(\bm{\theta}_{W})=\frac{1}{N}\sum_{i=1}^{N}\|\bm{z}_{t^{i}}(\bm{s}^{i})-\bm{u}_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})\|_{2}^{2}.$ (10)
Note, in the original representation of data-driven discovery, there is no measurement error associated with the observation in (10). The solution proceeds by optimizing the weighted combination
@@
where we now explicitly note that $J_{P}$ depends on the PDE parameters $\bm{\theta}_{m}$. After this optimization, one has learned the PDE parameters and the associated neural network can be used to obtain an estimate of $\bm{u}_{t}(\bm{s})$ at desired locations in space and time.
Importantly, “data-driven discovery of PDEs” does not mean finding a functional form of PDEs in this case, but rather indicates estimating parameters within PDEs. The use of statistical models and deep neural networks to learn the functional form of mechanistic models is a growing area of research, but beyond the scope of this article. For recent overviews, see North et al., (2023), Sahimi, (2024), and North et al., (2025).
@@ 4.2 Bayesian PINNs
The PINN approach described above does not explicitly account for uncertainty in the data, model, or parameters. It is quite straightforward to account for measurement error as in the variational approach presented earlier; we can modify the data portion of the loss function (10) by assuming that the measurement errors are Gaussian. For example, we might assume that we have observation models of a univariate process (although this can be generalized to a multivariate process) $z_{t^{i}}(\bm{s}^{i})=u_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})+\epsilon_{t^{i}}(\bm{s}^{i}),$ where $\epsilon_{t^{i}}(\bm{s}^{i})\sim\text{Gau}(0,\sigma^{2}_{u})$. Similarly, in principle we could have observations of $r_{t}(\bm{s})$ (the PDE residual term), boundary condition $B(\cdot)$ and initial conditions $g_{I}(\cdot)$ that also have independent Gaussian errors and associated error variances. All these distributions are conditioned on the solution of the PDE, but in the PINN context, that is completely determined given the NN and PDE model parameters, $\bm{\theta}_{W},\bm{\theta}_{m}$. Note, in principle, one could have dependent errors as in the variational and physical-statistical approach above.
It is then natural, following from the Bayesian NN literature (e.g., Neal,, 2012), to specify prior distributions on the NN weight and bias parameters, $\bm{\theta}_{W}$; known as a B-PINN (e.g., Yang et al.,, 2021). That is, assume we generically have conditional data distributions for the process, PDE residual, boundary and initial conditions given by $[D_{u}|\sigma^{2}_{u},\bm{\theta}_{W},\bm{\theta}_{m}]$, $[D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\bm{\theta}_{m}]$, $[D_{bc}|\sigma^{2}_{bc},\bm{\theta}_{W},\bm{\theta}_{m}]$ and $[D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W},\bm{\theta}_{m}]$, respectively. Each of these terms is implicitly conditioned on $u_{NN,t}$, but as mentioned above, since the NN is completely determined conditioned on its weights and biases, it is not necessary to write this explicitly. In addition, as in the case of PNMs, unless we have a forcing term in the PDE, we assume implicitly that “observations” of the residuals are equal to zero on average – that is, conditioned on the parameters, the relationship in (1) holds perfectly up to some arbitrary white noise (represented by $\sigma^{2}_{r}$). This has the effect of including a term equivalent to (9) in the un-normalized posterior representation (12) below. That is, we can either write this as a “data model” with known zero observations or as a regularization prior term. For the latter, we might write the prior distribution as:
$\frac{\partial{u_{NN}(\bm{\theta}_{W})}}{\partial{t}}\sim\text{Gau}(\mathcal{M}_{u}({u_{NN}(\bm{\theta}_{W})};\bm{\theta}_{m}),\sigma^{2}_{r}),$ (11)
@@
$[D_{u},D_{r},D_{bc},D_{ic}|\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}]=[D_{u}|\sigma^{2}_{u},\bm{\theta}_{W},\bm{\theta}_{m}][D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\bm{\theta}_{m}][D_{bc}|\sigma^{2}_{bc},\bm{\theta}_{W},\bm{\theta}_{m}][D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W},\bm{\theta}_{m}],$
@@
where $\bm{\theta}_{D}\equiv(\sigma^{2}_{u},\sigma^{2}_{r},\sigma^{2}_{bc},\sigma^{2}_{ic})^{\prime}$ and assume we have prior distributions $[\bm{\theta}_{W}]$, $[\bm{\theta}_{m}]$, and $[\bm{\theta}_{D}]$. Then, Bayes’ rule gives
$[\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}|D]\propto[D|\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}][\bm{\theta}_{W}][\bm{\theta}_{m}][\bm{\theta}_{D}],$ (12)
@@
where $D\equiv\{D_{u},D_{r},D_{bc},D_{ic}\}$. The intractable posterior (12) is typically sampled from using MCMC methods (e.g., Hamiltonian Monte Carlo, Langevin Dynamics), variational methods, or ensemble methods (see the comprehensive review in Psaros et al.,, 2023).
@@
A significant challenge with Bayesian NNs (and, hence, B-PINNs) is the specification of reasonable prior distributions for the NN parameters $\bm{\theta}_{W}$. In general, such parameters are not identifiable, which can make MCMC sampling challenging without extra constraints (e.g., McDermott & Wikle,, 2019), the specification of informative priors (e.g., through empirical Bayes or meta-learning; e.g., Meng et al., (2022)), or special classes of priors (e.g., Meng et al.,, 2022; Sell & Singh,, 2023; Castillo & Egels,, 2024).
@@
Note that the B-PINN framework presented here does not include an explicit term for model error, although $\sigma^{2}_{r}$ does account for minor discrepancies in the PDE. But, this does not respect the fact that the PDE dynamics are not usually sufficient to model the real world process. In other words, the real-world dynamical process is latent and is motivated by the PDE, but is typically more complicated. This is exactly the motivation for denoting a distinction between the true latent process, $u$, and the PDE solution $\tilde{u}$ in Section 2.2 and the NN solution $u_{NN}$ here. Such discrepancy terms are common in the computer model experiment literature (e.g., Gramacy,, 2020). In the context of B-PINNs, a discrepancy term can be added in the data model, e.g., $z_{t^{i}}(\bm{s}^{i})=u_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})+\delta_{t^{i}}(\bm{s}^{i};\bm{\theta}_{\delta})+\epsilon_{t^{i}}(\bm{s}^{i})$, where $\delta_{t^{i}}(\bm{s}^{i};\bm{\theta}_{\delta})$ is a discrepancy term that depends on parameters $\bm{\theta}_{\delta}$. In the statistical surrogate and mechanistically-motivated statistics literature, this discrepancy is typically modeled as a GP, a functional (basis) expansion, or a machine learning surrogate. In the B-PINN context, it is natural to model the discrepancy as an additional (separate) neural network (e.g., Zou et al.,, 2024), with the discrepancy used to correct for the misspecification of the PDE when observing PDE residuals. In either case, one must be careful to ensure that there is sufficient prior information regarding the spatial (and possibly temporal) scale of the discrepancy process to ensure identifiability of the solution (see our example in Section 6 below for more discussion).
@@
Although the B-PINN framework is an increasingly useful paradigm to account for uncertainty in PINN applications, there are still potential limitations in its lack of an explicit hierarchical formulation, lack of accommodation of dependent random effects, focus on Gaussian data distributions, and lack of more complicated measurement/sampling errors. In the next section, we show how the PINN fits into a more general deep mechanistically-informed BHM framework that can accommodate these concerns.
5 General Mechanistically-Informed Deep Dyamic Spatio-Temporal Model (GMID-DSTM)
Here, we present a general mechanistically-informed deep dynamic spatio-temporal model (GMID-DSTM) for statistical analysis of spatio-temporal data. The paradigm is flexible enough to include the examples presented previously (e.g., physical-statistical models, B-PINNs) as special cases. Importantly, this framework also can account for non-Gaussian data, model output as data, dependent random effects, and covariate effects. We describe the approach below following the spatio-temporal hierarchical modeling principles outlined in Cressie & Wikle, (2011).

### v2

@@ 3.2.2 Probabilistic Numerical Methods
Over the last decade, there has been growing interest in modernizing the ideas associated with treating differential equation solutions as random processes (e.g., larkin1972gaussian; diaconis1988bayesian). These ideas, which come under a broad set of methods called probabilistic numerical methods (PNMs), have provided a comprehensive toolkit to consider uncertainty quantification in the solution of differential equations (e.g., see hennig2015probabilistic; chkrebtii2016bayesian; cockayne2019bayesian, and references therein). These methods have the advantage of being rigorous in their convergence guarantees and are straightforward to implement.
Many PNM methods for solving PDEs with uncertainty assume the solution has a GP prior. Then observations of boundary conditions and forcing terms of the PDE at a finite set of (collocation) points act as data that allow for a posterior distribution over the PDE solution. For a simple illustration, in the space-only case, we might have the following GP prior:
@@
$f(s_{i})=\mathcal{M}(u(s_{i}))+\epsilon_{i},\;\;\epsilon_{i}\sim\text{N}(0,\sigma^{2}_{\epsilon}),$
where $\sigma^{2}_{\epsilon}$ controls the fidelity of the PDE solution. Given the prior and data models are normal, the posterior distribution of $u(s)$ given $f(s_{1}),\ldots,f(s_{m})$ are available in closed form via a normal distribution. In practice, one includes a boundary condition process and, for time-varying PDEs, an initial condition. In addition, one typically chooses the collocation points at which to obtain “observations” $f(s_{i})$ to facilitate computation. This approach is often embedded in a BHM with real world observations to learn the parameters of the underlying PDE (i.e., the inverse problem).
The strength of the PNM approach is its principled uncertainty quantification, which comes from the numerical approximation (e.g., the $f(s_{i})$ observations) and, in the case of inverse problems, real-world observations. The inherent regularization of the Bayesian prior also provides some robustness to ill-posedness. Perhaps most importantly, PNM’s strong connections to traditional numerical methods have been established for special cases (e.g., linear PDEs and GP priors). That said, these methods are much more difficult to implement for nonlinear PDEs, do not scale well for very high-dimensional problems, and have not been used as much in deep BHMs with multiple sources of process uncertainty, as in the physical-statistical paradigm. However, the Bayesian PINN framework described in the next section might be considered a neural network based PNM.
@@ 4 Physics-Informed Neural Networks (PINNs)
In the spirit of the Sasaki’s variational approach described above in Section 3.1, consider fitting a deep neural network (NN) but constraining the solution to follow a PDE. This is the essence of PINNs (Raissi2017; Raissi2017a; Raissi2018d; Raissi2019; Raissi2020). Such approaches can be used in the context of fitting a PDE in the presence of data (as in data assimilation), or solving a PDE explicitly (presumably for the sake of efficiency). We are interested here in the situation where we have data corresponding to the latent process of interest, initial conditions, or boundary conditions. In this section, we provide a brief overview of the PINN approach and its Bayesian counterpart (B-PINN).
@@ 4.1 PINN Formulation
PINNs, first proposed by Raissi2017; Raissi2017a, mainly consist of two components: (1) a neural network for prediction and (2) a physically-inspired equation (e.g., PDE) to impose constraints in the loss function. We denote the deep neural network solution to the PDE at time $t$ and location $\bm{s}$ as $\bm{v}_{NN,t}(\bm{s};\bm{\theta}_{W})$, where $\bm{\theta}_{W}$ represents the neural network parameters (weights and biases). Note, $\bm{v}_{NN,t}$ is an approximation of the PDE solution ${\bm{v}}_{t}$. Indeed, since this is not a perfect representation it is best to make the distinction as follows by defining a PDE “residual” (as in the PNM approach described in Section 3.2.2). Let $\bm{\theta}_{m}$ correspond to PDE parameters as in (1), and write
$\bm{r}_{t}(\bm{s};\bm{\theta}_{W},\bm{\theta}_{m})\equiv\frac{\partial{\bm{v}_{NN}(\bm{\theta}_{W})}}{\partial{t}}-\mathcal{M}({\bm{v}_{NN}(\bm{\theta}_{W})};\bm{\theta}_{m}),$ (8)
@@
$J_{IC}(\bm{\theta}_{W})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}\|{\bm{v}}_{NN,0}(\bm{s}_{IC}^{i};\bm{\theta}_{W})-\bm{g}_{I}(\bm{s}_{IC}^{i})\|_{2}^{2},$
$J_{BC}(\bm{\theta}_{W})=\frac{1}{N_{BC}}\sum_{i=1}^{N_{BC}}\|\bm{B}({\bm{v}}_{NN,t_{BC}^{i}}(\bm{s}_{BC}^{i};\bm{\theta}_{W}))\|_{2}^{2}.$
@@
where the specified non-negative constants $\{a_{IC},a_{BC},a_{P}\}$ allow one to focus more on different components of the solution if desired. The solution in this case does not require measurements of the process besides initial and boundary condition training data, so the PINN serves as an alternative to the traditional numerical solver (e.g., finite difference, finite element or spectral scheme). In this context, note a well-known limitation of the PINN approach is it can have difficulty in simulating multi-scale, chaotic, or turbulent behavior because the explicit forward (“causal”) nature of time is not considered (e.g., wang2022respecting). This limitation can be mediated in a number of ways, for example, by adding a time weighting function in the loss function (9). Regardless, some recent comparison studies have shown that traditional finite element solutions of PDEs work better and are as or more efficient than PINNs for most problems (e.g., grossmann2024can).
@@ 4.1.1 PINNs with Observational Data
$J_{D}(\bm{\theta}_{W})=\frac{1}{N}\sum_{i=1}^{N}\|\bm{z}_{t^{i}}(\bm{s}^{i})-\bm{v}_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})\|_{2}^{2}.$ (10)
Note, in the original presentation of data-driven discovery, there is no measurement error associated with the observation in (10). The solution proceeds by optimizing the weighted combination
@@
where we now explicitly note that $J_{P}$ depends on the PDE parameters $\bm{\theta}_{m}$. After this optimization, one has learned the PDE parameters and the associated neural network can be used to obtain an estimate of the latent process of interest, $\bm{u}_{t}(\bm{s})$, at desired locations in space and time.
Importantly, “data-driven discovery of PDEs” does not mean finding a functional form of PDEs in this case, but rather indicates estimating parameters within PDEs. The use of statistical models and deep neural networks to learn the functional form of mechanistic models is a growing area of research, but beyond the scope of this article. For recent overviews, see north2023review, sahimi2024physics, and north2025bayesian.
@@ 4.2 Bayesian PINNs
The PINN approach described above does not explicitly account for uncertainty in the data, model, or parameters. It is quite straightforward to account for measurement error as in the variational approach presented earlier; we can modify the data portion of the loss function (10) by assuming that the measurement errors are normal. For example, we might assume that we have observation models of a univariate process (although this can be generalized to a multivariate process) $z_{t^{i}}(\bm{s}^{i})=v_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})+\epsilon_{t^{i}}(\bm{s}^{i}),$ where $\epsilon_{t^{i}}(\bm{s}^{i})\sim\text{N}(0,\sigma^{2}_{z})$. Similarly, in principle we could have observations of $r_{t}(\bm{s})$ (the PDE residual term), boundary condition $B(\cdot)$ and initial conditions $g_{I}(\cdot)$ that also have independent normal errors and associated error variances. All these distributions are conditioned on the solution of the PDE and PDE model parameters, which in the PINN context are completely determined given the NN and PDE model parameters, $\bm{\theta}_{W},\bm{\theta}_{m}$. Note, in principle, one could have dependent errors as described in the variational and physical-statistical approaches above.
It is then natural, following from the Bayesian NN literature (e.g., neal2012bayesian), to specify prior distributions on the NN weight and bias parameters, $\bm{\theta}_{W}$; known as a B-PINN (e.g., yang2021b). That is, assume we generically have conditional data distributions for the process, PDE residual, boundary and initial conditions given by $D_{z}|\sigma^{2}_{z},\bm{\theta}_{W}]$, $D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\bm{\theta}_{m}]$, $D_{bc}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ and $D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W}]$, respectively. Note, we use $D_{z}$ to represent the observations $\{z_{t^{i}}(\bm{s}^{i})\}$ described above to provide a more general and consistent notation for the observations. Each of these terms is implicitly conditioned on $v_{NN,t}$, but as mentioned above, since the NN is completely determined conditioned on its weights and biases, it is not necessary to write this explicitly. In addition, as in the case of PNMs, unless we have a forcing term in the PDE, we assume implicitly that “observations” of the residuals are equal to zero on average – that is, conditioned on the parameters, the relationship in (1) holds perfectly up to some arbitrary white noise (represented by $\sigma^{2}_{r}$). This has the effect of including a term equivalent to (9) in the un-normalized posterior representation (12) below. That is, we can either write this as a “pseudo data model” with known zero observations or as a regularization prior term. For the latter, we might write the prior distribution as:
$\frac{\partial{v_{NN}(\bm{\theta}_{W})}}{\partial{t}}\sim\text{N}(\mathcal{M}({v_{NN}(\bm{\theta}_{W})};\bm{\theta}_{m}),\sigma^{2}_{r}),$ (11)
@@
$D_{z},D_{r},D_{bc},D_{ic}|\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}]=[D_{z}|\sigma^{2}_{z},\bm{\theta}_{W}][D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\bm{\theta}_{m}][D_{bc}|\sigma^{2}_{bc},\bm{\theta}_{W}][D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W}],$
@@
where $\bm{\theta}_{D}\equiv(\sigma^{2}_{z},\sigma^{2}_{r},\sigma^{2}_{bc},\sigma^{2}_{ic})^{\prime}$. The parameter $\bm{\theta_{m}}$ is only included for the conditional distribution of $D_{r}$ because PDE residuals are the only quantities that depend on the physics parameters; the other quantities’ conditional distributions are only dependent on the neural network, and consequently only $\bm{\theta}_{W}$. Now, assume we have prior distributions $\bm{\theta}_{W}]$, $\bm{\theta}_{m}]$, and $\bm{\theta}_{D}]$. Then, Bayes’ rule gives
$\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}|D]\propto[D|\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}][\bm{\theta}_{W}][\bm{\theta}_{m}][\bm{\theta}_{D}],$ (12)
@@
where $D\equiv\{D_{z},D_{r},D_{bc},D_{ic}\}$. The intractable posterior (12) is typically sampled from using MCMC methods (e.g., Hamiltonian Monte Carlo, Langevin Dynamics), variational methods, or ensemble methods (see the comprehensive review in psaros2023uncertainty).
@@
A significant challenge with Bayesian NNs (and, hence, B-PINNs) is the specification of reasonable prior distributions for the NN parameters $\bm{\theta}_{W}$. In general, such parameters are not identifiable, which can make MCMC sampling challenging without extra constraints (e.g., mcdermott2019bayesian), the specification of informative priors (e.g., through empirical Bayes or meta-learning; e.g., meng2022learning), or special classes of priors (e.g., meng2022learning; sell2023trace; castillo2024posterior).
@@
Note that the B-PINN framework presented here does not include an explicit term for model error, although $\sigma^{2}_{r}$ does account for minor discrepancies in the PDE. But, this does not respect the fact that the PDE dynamics are not usually sufficient to model the real world process. In other words, the real-world dynamical process is latent and is motivated by the PDE, but is typically more complicated. This is exactly the motivation for denoting a distinction between the true latent process, $u$, and the PDE solution $v$ in Section 2.2 and the NN solution $v_{NN}$ here. Such discrepancy terms are common in the computer model experiment literature (e.g., gramacy2020surrogates). In the context of B-PINNs, a discrepancy term can be added in the data model, e.g., $z_{t^{i}}(\bm{s}^{i})=v_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})+\delta_{t^{i}}(\bm{s}^{i};\bm{\theta}_{\delta})+\epsilon_{t^{i}}(\bm{s}^{i})$, where $\delta_{t^{i}}(\bm{s}^{i};\bm{\theta}_{\delta})$ is a discrepancy term that depends on parameters $\bm{\theta}_{\delta}$. In the statistical surrogate and mechanistically-motivated statistics literature, this discrepancy is typically modeled as a GP, a functional (basis) expansion, or a machine learning surrogate. In the B-PINN context, it is natural to model the discrepancy as an additional (separate) neural network (e.g., zou2024correcting), with the discrepancy used to correct for the misspecification of the PDE when observing PDE residuals. In either case, one must be careful to ensure that there is sufficient prior information regarding the spatial (and possibly temporal) scale of the discrepancy process to ensure identifiability of the solution (see our example in Section 6 below for more discussion).
@@
Although the B-PINN framework is an increasingly useful paradigm to account for uncertainty in PINN applications, there are still potential limitations in its lack of an explicit hierarchical formulation, lack of consideration of covariate effects, lack of accommodation of dependent random effects, focus on normal data distributions, and lack of more complicated measurement/sampling errors. In the next section, we show how the B-PINN fits into a more general deep BHM framework for dynamic spatio-temporal processes that can accommodate these concerns.
5 Bayesian Hierarchical Model PINNs (BHM-PINNs)
A strength of an overview is drawing generalizations and connections to a larger body of work. As part of this overview of PINNs, we show here that the B-PINN (yang2021b) fits naturally into the BHM framework that has long been used to model dynamic spatio-temporal processes in statistics (e.g., see cressie2011statistics, for a review). It is important to note that this is not a new methodology, but simply a new use case of the spatio-temporal BHM framework. We denote this a BHM-PINN and show it is a flexible approach to bring mechanistic-inspired PDE dynamics to spatio-temporal modeling in a paradigm that can account for non-normal data distributions, covariate effects, and residual spatial or spatio-temporal processes. This is analogous to how mechanistic-informed models are included in physical-statistical BHMs. Thus, the BHM-PINN described in the following subsections generalizes both the traditional BHM spatio-temporal physical-statistical modeling framework and the PINN modeling framework.